find the equations of the asymptotes of the hyperbola x^2 -9y^2=18
Answer by lwsshak3(11628)

find the equations of the asymptotes of the hyperbola
x^2 -9y^2=18
x^2/18-y^2/2
hyperbola has a horizontal transverse axis with center at the origin.
Its standard form of equation: x^2/a^2-y^2/b^2
For given hyperbola:
a^2=18
a=√18
b^2=2
b=√2
For hyperbolas with a horizontal transverse axis:
slopes of asymptotes=±b/a=±√(2/18)=±(1/3)
2 equations for asymptotes:
y=x/3
y=-x/3